the perimeter fence of a rectangle community is 310m. Its length is 10 meter less than twice its width. Find the dimensions of the park.
Answer by ewatrrr(24785)


Hi,
Let x represent the width, then (2x-10m) would be the length
2*(2x - 10m) + 2*x = 310m (the perimeter)
solving for x
4x - 20m + 2x = 310m
6x = 330m
x = 55m, the width. the length is 100m (2*55m - 10)
checking our answer
110 + 200 = 310